OpenStudy (anonymous):

3sqrt16 * 3sqrt4 the 3 is above the sqrt

OpenStudy (bahrom7893):

what does that mean?

OpenStudy (bahrom7893):

How can you rewrite that expression?

OpenStudy (anonymous):

I don't know. I need help solving it. It just says 3sqrt16 times 3sqrt 4 and it looks like the 3 is above the sqrt and it's smaller

OpenStudy (bahrom7893):

What is the topic in the textbook? You need to have at least some basic knowledge

OpenStudy (bahrom7893):

We're not here to do the homework.

OpenStudy (bahrom7893):

We're here to teach, and you need to be willing to put in the work.

OpenStudy (anonymous):

Do you have any directions? o.o

OpenStudy (anonymous):

I don't know how to do it. I gave you what it says and I can't upload a picture. I need help, so you can't get upset with me for not knowing how to do it and get frustrated when I ask how?

OpenStudy (anonymous):

\[\sqrt[3]{16}\times \sqrt[3]{4}=\left( 16 \right)^{\frac{ 1 }{ 3 }} \times 4^{\frac{ 1 }{ 3 }}=\left( 4^2 \right)^{\frac{ 1 }{ 3 }} \times 4^{\frac{ 1 }{ 3 }} =?\]

OpenStudy (anonymous):

16x^4/9

OpenStudy (anonymous):

\[4^{2 \times \frac{ 1 }{ 3 }} \times 4^{\frac{ 1 }{ 3 }}=4^{\frac{ 2 }{ 3 }+\frac{ 1 }{ 3 }}=4^{\frac{ 2+1 }{ 3 }}=4^{\frac{ 3 }{ 3 }}=4^1=4\]